### v1

# The Failed Migration of Academic Twitter

Abstract
Following change in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some. Our study looks at the dynamics of this migration. Utilizing publicly available user account data, we track the posting activity of academics on Mastodon over a one year period. Our analyses reveal significant challenges sustaining user engagement on Mastodon due to its decentralized structure as well as competition from other platforms such as Bluesky and Threads. The movement lost momentum after an initial surge of enthusiasm as most users did not maintain their activity levels, and those who did faced lower levels of engagement compared to Twitter. Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focusing on migrating social connections for long-term user engagement.
Keywords:
Social Media Mastodon Twitter Social Movement Computational Social Science Decentralized Platforms.
**footnotetext: These authors contributed equally to this work.
1 Introduction
Social media platforms have become essential outlets for communication, information dissemination, and professional networking. Among these platforms, Twitter (now known as “X”) has gained significant popularity, often referred to as virtual town square for its role facilitating public discourse [10]. Academics and journalists populate sub-communities on Twitter and use the platform for professional communication [7, 9]. Academics have extensively used Twitter to share research findings, announce academic events such as workshops and conferences, and contribute to wider scholarly discourse [11, 8]. However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives [2].
Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their topics of interest. Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network [14]. This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control [12]. Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels [12].
The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities. Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform[13]. Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process [3].
Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence [5]. Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities [2]. Analysis showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms [4]. A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture [6].
Despite prior work, there remains a gap in the research concerning specific professional communities, as previous work has focused on the general population of social media platform users. In this study, we address this gap by examining the migration patterns and behaviors of academic users, providing a more nuanced understanding of how these particular communities navigate platform transitions. Our study addresses the following research questions:
RQ1: To what extent were academics engaged in the migration from Twitter to Mastodon, and how can the success or failure of this migration be assessed?
RQ2: Which academic disciplines have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
Our findings indicate that majority of the academic researchers moved to Mastodon failed to maintain their posting activity and follower engagement. These results underscore the importance of not only transferring accounts but also migrating social connections to ensure the success of social media transitions.
2 Data
We utilized a public Github repository containing lists of academics on Mastodon [^1] https://github.com/nathanlesage/academics-on-mastodon?tab=readme-ov-file. These lists were curated by researchers from 50 distinct disciplines, who volunteered to be added to the lists. Each list corresponds to a specific research discipline, containing account details of researchers within that research area. If a list for a particular discipline does not exist, one can start a new list and provide a corresponding form. This form allows researchers to add themselves to the relevant list based on the discipline they identify with. While the database continues to expand, for the purpose of tracking, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022. In total, we tracked the Mastodon profiles and public metrics of 7542 academic users using a customized Mastodon API. Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of the user activity and engagement over the span of a year.
For cross-platform analysis, we also collected tweets via Zeeschuimer [^2] https://github.com/digitalmethodsinitiative/zeeschuimer using the query "mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01". Since many tweets were included because the username contained the word "mastodon" (as some users use their Mastodon account name as their Twitter username, we further filtered out these tweets). We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets.
3 Results
3.1 Attrition of academic accounts on Mastodon
Figure 1 displays a count heat map illustrating the relationship between “last status time" and “record time”, aggregated by month. The “last status time” refers to the time of the user’s most recent post, while “record time" indicates the time at which the data was collected.
In the initial month of data collection (November 2022), there were 7,505 active scholars who had posted within that same month. This number experienced a gradual decline over the subsequent seven months, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023. The figure also shows that 1,237 scholars were active in November 2022 but then not again throughout the remainder of our data collection. We observe a significant rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
Figure 1: Heat map of user activity. Last status vs. record time (aggregated by month).
3.2 Engagement with academic accounts on Mastodon
Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. To do, so we first characterize accounts based on their activity on the platform.
3.2.1 Active users.
We define active users as scholars who engaged on Mastodon during the initial Twitter-to-Mastodon migration, which began in late October 2022. For our purposes, we’ve defined this initial period between October 15 and November 30, 2022. We further divide active users into sub-categories based on level of engagement: one-time visitors; short-term, long-term, and persistent adopters. Each category reflects the duration of their activity on Mastodon following the migration.
One-time visitors. These academics created an account on Mastodon at the beginning of the Twitter-to-Mastodon migration, posted only within the first week of joining and then abandoned the platform ($n_{days}\leq 7$). This brief engagement suggests a possible exploration of the platform following the migration advocacy, without finding enough value or community to continue their activity.
Short-term adopters. This group includes those who actively used Mastodon for more than one week but did not persist beyond 100 days. They left the platform possibly due to unmet expectations or a stronger pull back to more familiar platforms like Twitter.
Long-term adopters. These users remained active on Mastodon from the time of migration for relatively long time ($100<n_{days}\leq 300$) but have been inactive.
Persistent adopters. These users remained active on Mastodon from initial migration though the end of our analyses ($n_{days}>300$). Their sustained engagement indicates successful integration into the Mastodon community.
3.2.2 Inactive users.
Academics in this category were already using Mastodon before the migration wave from Twitter. Despite their early presence on the platform, they remained inactive during the migration period. One user was inactive during the first month but engaged briefly afterward. Inactive users are not included in subsequent analyses. The distribution of users is given in Table 1. Note that only 35.4% of academic users in our dataset are active and consistent in their use of Mastodon. The majority demonstrate sporadic engagement.
Category Subcategory Count
Active Users One-time visitors 1,183
Short-term adopters 1,624
Long-term adopters 2,065
Persistent adopters 2,665
Inactive Users One-Time Visitors 4
Returned Users 1
Table 1: Number of users, by activity.
We calculate the average following, follower and post counts for each user category, aggregated by month show in Figure 2. Our analysis reveals a sharp increase in all three public metrics at the onset of the migration period, specifically from November 2022 to December 2022. However, beyond this period, the rate of increase for following and follower metrics slows considerably. For one-time, short-term, and long-term users, growth in following and follower counts stagnated beginning in early 2023. As these users were not regularly contributing content, their visibility and potential to attract followers were likewise limited. In contrast, persistent users continued to post regularly. Despite their consistent activity, however, growth in their connections remained slow. This slow growth highlights the challenges faced in establishing and expanding social ties and communities on Mastodon.
Figure 2: Average public metrics and statuses count for active users.
3.3 Field-specific activity for academic accounts on Mastodon
As discussed in Section 2, users self-identified to research disciplines in public lists. We further categorize these disciplines into 9 fields, consistent with existing taxonomies of academic disciplines [1]. Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers. Inactive users were excluded from subsequent analyses. Disciplinary statistics of the 7,537 academics in our datasets are provided in Figure A5 in the Appendix.
We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table 2). The highest proportion of persistent users is observed within multidisciplinary researchers. The public metrics aggregated by field and by month are shown in Figure A4 in the Appendix. Researchers in the physical sciences and mathematics have notably higher average number of posts, indicating their consistent and significant contribution to discussions on Mastodon. Scholars in medicine and health sciences exhibit relatively less frequent posting.
Field One Time (%) Short Term (%) Long Term (%) Persistent (%)
Social and Behavioral Sciences 15.7 21.4 27.9 35.0
Arts and Humanities 20.5 19.4 24.8 35.3
Physical Sciences and Mathematics 12.0 21.2 27.8 39.0
Life Sciences 10.7 20.2 41.7 27.4
Education 21.3 26.2 24.2 28.3
Law 17.9 28.3 26.6 27.3
Medicine and Health Sciences 10.8 30.3 29.3 29.6
Multidisciplinary 10.8 18.3 28.8 42.1
Table 2: Distribution of user engagement by academic field.
3.4 Twitter and Mastodon cross-platform analysis
To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 2). We plot the daily frequency of these tweets over time in Figure 3. Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement.
Conversations reflect difficulties and concerns during the transition. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
"Looks like Mastodon isn’t an option anyway. Academic servers aren’t taking new accounts at the moment."
"I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am <scholar’s Mastodon account>."
"As you know, I’m over on Mastodon too. Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does."
However, this initial surge in discussion was short-lived. The frequency of related tweets sharply declined.
Figure 3: Dynamic frequency of Twitter conversations related to academics migrating to Mastodon.
Users expressed frustration over the platform’s lack of centralized theme and purpose, and its fragmented community structure.
"Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity. but the *big* servers that most ppl are on are… kinda academic? kinda techy? very… not fun? it feels like linkedin or facebook groups to a lot of people."
"My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity."
Despite these challenges, ongoing efforts to promote Mastodon indicate a persistent belief in its potential. For instance, in June 2023, one scholar posted,
"Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd"
Amongst the discussions, we observe concurrent mentions of other platforms such as Bluesky [^3] https://bsky.app/ and Threads [^4] https://www.threads.net/. These have diluted the migration as users explore various platforms for different purposes. For instance,
"I’m curious re categories of people I follow and where they go. IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?)."
"Hey if anyone’s still here on science twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic"
3.4.1 Case Study: Physical Sciences and Mathematics Scholars.
We observe that scholars in the physical sciences and mathematics were amongst the most engaged on Mastodon. We randomly sampled 100 scholars in these fields who had provided Twitter account information and matched their Mastodon accounts with their Twitter accounts. We recorded whether they are still active on Twitter, and whether they still provide their Mastodon account information in their username, profile description, or pinned tweets. Additionally, we collected follower and following statistics of the sampled scholars on Twitter, all statistics shown in Table 3. We observe that 86% of sampled scholars are still active on Twitter, and among these active users, 62.8% do not provide Mastodon information in easily accessible places, e.g., username, bio, pinned tweets, indicating no signs of migration or intention to migrate. These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform. The rest of the active users still make their Mastodon information accessible to the community. For instance, one scholar noted:
"Hi ya’ll for those moving over to #Mastodon you can currently find me <Scholar’s Mastodon Account>."
This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon. In contrast, 5% of the sampled scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter. For example, one scholar’s last post reads,
"I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!"
Some, 9% of our sampled scholars, chose to lock their accounts as a sign of leaving Twitter, leaving a message in their profile description directing followers to their Mastodon accounts.
Contain Mastodon Info Not Contain Mastodon Info Following Count Follower Count
Active 32 54 Mean: 1593, std: 2684 Mean:11022, std: 30146
Inactive 4 1 Mean: 330, std: 451 Mean: 848, std: 1396
Locked 7 2 Mean: 1551, std:1361 Mean: 1631, std: 1837
Table 3: Engagement statistics of 100 sampled scholars.
In addition to segmentation, many scholars direct their followers to Threads and Bluesky instead of Mastodon, with 2% providing Threads information and 6% providing Bluesky information among the sampled scholars. This trend indicates that Mastodon faces significant competition from other emerging platforms, making it even more challenging for Mastodon to become a preferred alternative for academic discourse.
In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously. Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community. Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4. Our analysis reveals that 55% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users). Among those who are inactive on Twitter or have locked their accounts, 71% are long-term or persistent users of Mastodon. This indicates that while Mastodon has some loyal users, their numbers are relatively small compared to the entire academic community.
One_Time Short_Term Long_Term Persistent
Active 12 19 28 27
Inactive 1 1 2 1
protected 0 2 1 6
Table 4: Mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
4 Discussion and Conclusion
Our analysis of professional communities migrating between social media platforms highlights the unique challenges these users face during such transitions. While decentralized platform offered a promising alternative, the established communities and clear audience engagement on Twitter proved too significant to overcome. The fragmented communities on Mastodon lacked the cohesiveness found on Twitter. Our results show that the initial surge in Mastodon sign-ups did not translate into sustained user engagement.
Among the users we tracked, the majority failed to maintain the same level of activity. This could be attributed to steep learning curve compared to twitter as well as competition from more user-friendly platforms like Bluesky and Threads. These alternatives might have drawn away potential users exacerbating Mastodonś difficulties in becoming a central hub for professionals communities. Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter which is crucial for sustaining professional discourse.
Our study highlights that without a good alternative, professionals will naturally gravitate back to familiar platforms where their communities are already established through years. Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement. While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions. Successful transitioning involves creating communities that rival established social media networks and user experiences of platforms such as Twitter.
{credits}
[…]
There is no competing interest.