### v2

# The Failed Migration of Academic Twitter

Abstract.
Following changes in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some. Our study looks at the dynamics of this migration. Utilizing publicly available user account data, we track the posting activity of academics on Mastodon over a one year period. We also gathered follower-followee relationships to map internal networks, finding that the subset of academics who engaged in migration were well-connected. However, this strong internal connectivity was insufficient to prevent users from returning to Twitter/X. Our analyses reveal significant challenges sustaining user engagement on Mastodon due to its decentralized structure as well as competition from other platforms such as Bluesky and Threads. The movement lost momentum after an initial surge of enthusiasm where the main network was fully established as most users did not maintain their activity levels, and those who did faced lower levels of engagement. Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focus on community building for long-term user engagement.
Social Media, Mastodon, Twitter, Social Movement, Computational Social Science, Decentralized Platforms
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Social networking sitesccs: Human-centered computing Social mediaccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Social networks
1. Introduction
Social media platforms have become essential outlets for communication, information dissemination, and professional networking. Among these platforms, Twitter (now known as “X”) has gained significant popularity and has been likened to a virtual town square for its role facilitating public discourse (Mascaro and Goggins, 2012). Academics and journalists populate sub-communities on Twitter and use the platform for professional communication (Jurkowitz and Gottfried, 2022; Malik et al., 2019). Academics have extensively used Twitter to share research findings, announce academic events such as conferences, and contribute to wider scholarly discourse (Mohammadi et al., 2018; Ebner, 2013; Mahrt et al., 2014). However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives (Cava et al., 2023).
Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their areas of interest. Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network (Zignani et al., 2018). This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control (Raman et al., 2019). Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels (Raman et al., 2019).
The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities. Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014). Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts of change (Fiesler and Dym, 2020).
Additionally, research has shown that user migration is influenced by perceived platform attractiveness and peer effects (Hou and Shiau, 2020). A study of users migrating from Twitter to Mastodon post-Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023). Analyses showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023). A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
Despite prior work characterizing the general population of platform users, there remains a gap in targeted research concerning specific professional communities. In this study, we address this gap by examining the migration patterns and behaviors of academic users, providing a more nuanced understanding of how these particular communities navigate platform transitions. Given the significant role that Twitter plays in an academic’s professional life (Knight and Kaye, 2016; Klar et al., 2020; Stewart, 2015), academics represent a uniquely valuable group for studying the migratory dynamics. Our study addresses the following research questions:
RQ1: To what extent were academics engaged in the migration from Twitter to Mastodon, and how can the success or failure of this migration be assessed?
RQ2: Which academic fields have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
Our findings indicate that majority of academic researchers who moved to Mastodon failed to maintain their posting activity and follower engagement over long term. Although migrated academic accounts formed strongly-connected internal networks, these connections were insufficient to sustain user engagement over time. Of note, however, communities that established field specific server instances had higher user retention rate compared to those on general-purpose servers. Cross platform analysis of most recent Twitter activity showed that most users returned to Twitter. These results underscore the challenges involved and the potential opportunities to build and foster sustainable communities at scale on new platforms.
2. Related Work
2.1. Characteristics of Mastodon
The decentralized structure Mastodon and similar platforms leads to unique user engagement dynamics that are different from centralized platforms. Research suggests that users are drawn to Mastodon for the ease of protecting their information from data mining and the absence of algorithms shaping their feeds (Lee and Wang, 2023). However, this can also make content discovery more difficult (Zignani et al., 2019). The decentralized structure encourages strong, tight-knit communities within individual server instances but fewer connections across instances (Zignani et al., 2018). These characteristics pose challenges for migration as well.
2.2. Cross-platform migration
Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014). The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities’ retention.
Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process (Fiesler and Dym, 2020). Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence (Hou and Shiau, 2020). Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023). Analysis showed Twitter to Mastodon migration is characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023). A similar study on Twitter-to-Mastodon migration reveals lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
Our work looks more deeply at the specific case of the relatively well-synchronized and coordinated movement of communities of academic users across platforms.
3. Data
Note on terminology: Twitter rebranded to X in July 2023, which occurred midway through our data collection. For simplicity, we use the terminology Twitter and tweets throughout the paper.
Our analyses are primarily based on a public Github repository containing lists of academics on Mastodon [^1] https://github.com/nathanlesage/academics-on-mastodon. These lists were curated by researchers from 50 distinct disciplines, who voluntarily added their own names and accounts. Each list corresponds to a research discipline, containing account details of researchers within that area, their Mastodon IDs, and a short description of their research. Some lists also included each user’s Twitter handle. While the database continues to expand, for the purpose of longitudinal analysis, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022. That is, the users in our dataset were part of a coordinated, synchronized migration effort in November 2022.
In total, we collected the Mastodon profiles and corresponding public metrics, e.g., number of followers, following, timestamps of most recent status of 7,542 academic user accounts using a customized Mastodon API [^2] https://github.com/halcy/Mastodon.py.git. Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of user activity and engagement over the span of a year. In addition, we collected the most recent Mastodon data related to follower and following accounts for each scholar to conduct network analyses based on fields of study and associated servers.
For cross-platform analyses, we used self-reported Twitter handles to gather scholars’ profile information. Of 7,542 Mastodon scholars, a subset of 3,131 scholars accurately provided Twitter handles suitable for data collection. Using the Twitter (“X”) Basic API, we retrieved profile information, most recent tweet IDs, and engagement metrics. Additionally, we used the Twitter ID collection endpoint to extract the posting time of their most recent tweet.
Finally, to understand discussions surrounding Mastodon in academic context, we collected tweets during the one year window via Zeeschuimer [^3] https://github.com/digitalmethodsinitiative/zeeschuimer using the boolean search “mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01”. Since many tweets were included because the username contained the word “mastodon” (as some users use their Mastodon account name as their Twitter username) we further filtered out these tweets. We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets.
Figure 1. Waterfall plot showing number of monthly active users (active=at least one post per month).
4. Results
Following, we present findings from multiple analyses conducted to understand the dynamics of the migration. First, we used the longitudinal data from Mastodon to understand user attrition and engagement trends over the year. Then, we used network analysis to understand the connectivity and distribution of academics across Mastodon servers. Next, we did a cross-platform analysis using the subset of users who provided Twitter handles comparing their activity patterns on Twitter and Mastodon. Finally, we ran exploratory analyses on Twitter posts related to the migration to understand user attitudes towards migration.
4.1. Longitudinal Analysis
4.1.1. Attrition of academic accounts on Mastodon
Figure 1 gives a count waterfall plot illustrating the change in number of active users aggregated by month and a line plot showing the monthly retention rate based on the previous month. In this context, we considered a user to be active if they made at least a post in a given month. In the initial month of data collection (November 2022), there were 7,505 active scholars who had posted within that same month. [^4] There were 1, 4, and 32 scholars had their last activity in April, May, and October 2022, respectively, which fell outside our collection window. These scholars were excluded from the waterfall plot but included in subsequent analysis. Over the next seven months, the number of active scholars experienced a decline, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023. We observe a subtantial rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
4.1.2. Engagement with academic accounts on Mastodon
Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. To do so, we first characterize accounts based on their activity.
Active users. We define active users as scholars who engaged on Mastodon during the initial Twitter-to-Mastodon migration, which began in late October 2022. For our purposes, we’ve defined this initial period between October 15 and November 30, 2022. We further divide active users into sub-categories based on level of engagement: one-time visitors; short-term, long-term, and persistent adopters. Each category reflects the duration of their activity on Mastodon following the migration.
One-time visitors. These academics created an account on Mastodon at the beginning of the Twitter-to-Mastodon migration, posted only within the first week of joining and then abandoned the platform ($n_{days}\leq 7$). This brief engagement suggests a possible exploration of the platform following the migration advocacy, without finding enough value or community to continue their activity.
Short-term adopters. This group includes those who actively used Mastodon for more than one week but did not persist beyond 100 days. They left the platform possibly due to unmet expectations or a stronger pull back to more familiar platforms like Twitter.
Long-term adopters. These users remained active on Mastodon from the time of migration for relatively long time ($100<n_{days}\leq 300$) but have been inactive.
Persistent adopters. These users remained active on Mastodon from initial migration though the end of our analyses ($n_{days}>300$). Their sustained engagement indicates successful integration into the Mastodon community.
4.1.3. Inactive users.
Academics in this category were already using Mastodon before the migration wave from Twitter. Despite their early presence on the platform, they remained inactive during the migration period. One user was inactive during the first month but engaged briefly afterward. Inactive users are not included in subsequent analyses. The distribution of users is given in Table 1. Note that only 35.4% of academic users in our dataset are active and consistent in their use of Mastodon. The majority demonstrate sporadic engagement.
User Category Subcategory Count
Active One-time visitors 1,183
Short-term adopters 1,624
Long-term adopters 2,065
Persistent adopters 2,665
Inactive One-Time Visitors 4
Returned Users 1
Table 1. Number of users, by activity.
We calculate the average following, follower and post counts for each user category, aggregated by month show in Figure 2. Our analysis reveals a sharp increase in all three public metrics at the onset of the migration period, specifically from November 2022 to December 2022. However, beyond this period, the rate of increase for following and follower metrics slows considerably. For one-time, short-term, and long-term users, growth in following and follower counts stagnated beginning in early 2023. As these users were not regularly contributing content, their visibility and potential to attract followers were likewise limited. In contrast, persistent users continued to post regularly. Despite their consistent activity, however, growth in their connections remained slow. This slow growth highlights the challenges faced in establishing and expanding social ties and communities on Mastodon. The persistent users’ continued efforts to engage and post suggest a commitment to the platform, yet the limited increase in followers indicates potential barriers in user retention and community building.
Figure 2. Average public metrics and statuses count for active users.
4.1.4. Field-specific activity differences among academic accounts on Mastodon
As discussed in Section 3, users self-identified their research disciplines in public lists. We aggregate these disciplines into 9 broad fields, consistent with existing taxonomies of academic disciplines (bepress., 2019). Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers. Inactive users were excluded from subsequent analyses. Disciplinary statistics of the 7,537 academics in our datasets are provided in Table A1 in the Appendix.
We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table 2). The highest proportion of persistent users is observed within multidisciplinary researchers. The public metrics aggregated by field and by month are shown in Figure A1 in the Appendix. Researchers in the physical sciences and mathematics have notably higher average number of posts, indicating their consistent and significant contribution to discussions on Mastodon. Scholars in medicine and health sciences exhibit relatively less frequent posting.
Field One Time (%) Short Term (%) Long Term (%) Persistent (%) Total Percent In-Degree Out-Degree
Social & Behavioral Sciences 15.7 21.4 27.9 35.0 2918 38.72 98881 91958
Arts & Humanities 20.5 19.4 24.8 35.3 1522 20.19 38692 39986
Physical Sciences & Mathematics 12.0 21.2 27.8 39.0 851 11.29 11453 11980
Life Sciences 10.7 20.2 41.7 27.4 84 1.11 338 648
Education 21.3 26.2 24.2 28.3 244 3.24 2249 4171
Law 17.9 28.3 26.6 27.3 587 7.79 21664 27450
Medicine & Health Sciences 10.8 30.3 29.3 29.6 287 3.81 4292 4923
Multidisciplinary 10.8 18.3 28.8 42.1 1040 13.80 56371 52818
Business 25 25 25 25 4 0.05 60 66
Table 2. Distribution of user engagement and node degree statistics by academic field.
4.2. Mastodon Network Analyses
To understand the connectivity within and between different fields, we collected the complete follower-followee data of all academic accounts on Mastodon, and constructed an internal connection network within each academic group, as well as a network aggregated by field of study.
4.2.1. Internal Network Aggregated by Field
Figure 3 shows the network connections across different fields. In this network, node size represents the number of academics in each field, while edge width indicates the frequency of connections between the user accounts from different fields. Given the nature of the field, certain fields such as Social and Behavioral Sciences are strongly connected to Multi-disciplinary field. Our analysis reveals that the academic migration has led to the formation of a well-connected internal network, suggesting that researchers actively attempted to establish a new community on Mastodon. However, despite these coordinated efforts, the migration still faced significant challenges, as reflected through the low retention rates shown in Section 4.1.
4.2.2. Internal Networks aggregated by Server
Prior studies have shown that the Mastodon’s server instance based structure makes it easier for users to discover and connect with others within the same instance than outside instance. This leads to stronger within instance ties compared to weaker links across instances (Zignani et al., 2019). Therefore, we analyzed the follower-followee network aggregated by Mastodon instances, shown in Figure 4.
The academic users were spread across a total of 342 servers, with 9 server instances being isolates without any connecting edges to other instances, which were removed for clarity. Table A2 in the Appendix shows network centrality metrics for each server and the distribution of the users across different servers. It also includes the official description of the server instances, indicating whether they are customized to a specific group of scholars. During the migration, different disciplines used different strategies: some used general-purpose servers like mastodon.social which host a wide variety of content beyond just academic content, while others took a coordinated approach on setting up their field specific server instances such as infosec.exchange and journa.host. Figure 4 shows the follower-followee connections across different servers showing certain general purpose instances such as mastodon.social are popular among the tracked academic users (containing 21.3% of the total) than other general purpose servers like mas.to.
Figure 3. internal field network (size: number of scholars in each field; edge weights: frequency of connections; edge color: same as the source).
4.2.3. Case study: Mastodon Community of Information Security
We also examined a discipline-specific server, unique to decentralized social platforms, and its user retention rate using a case study of the Information Security discipline. We chose this discipline based on our field-level engagement analysis shown in Figure A1 in Appendix, in the Appendix, which indicates that scholars in physical sciences and mathematics were among the most active on Mastodon. To further understand the impact of server instance choice on community formation, we visualized their internal network, highlighting the top five servers where these scholars are based. Figure A2 shows the internal follower-followee network for physical science and mathematics field, highlighting the top five servers from a total of 33 servers that these scholars use. Interestingly, scholars from Information Security formed a strongly connected community on a custom discipline specific server, unlike other disciplines that used general servers such as mastodon.social or fediscience.org.
To explore how a discipline specific server could influence user retention and engagement, we isolated the internal network of accounts within Information Security discipline. Figure 5 shows the internal network of the discipline, where most scholars congregated on a single server: infosec.exchange (highlighted in red). We compare the retention roles of scholars from the discipline with those from a discipline where users were concentrated in a general-purpose server and other sub-fields of similar size. We see a significantly high percentage of persistent users for information security scholars (56.8%) in comparison to other comparable disciplines. This suggests that establishing field- or discipline-specific communities or servers could foster a stronger sense of belonging, which may help retain scholars who have migrated, reducing the likelihood of their return to Twitter or seeking alternative platforms.
Figure 4. internal server network (removed isolated servers, red are servers with more than 100 scholars).
Figure 5. internal follower-followee network for Information Security community (removed isolates).
Category Info. Sec. Education Communication & Media Studies
Count Percent Count Percent Count Percent
One-Time 23 8.7 52 21.3 41 15.5
Short-Term 31 11.7 64 26.2 53 20.0
Long-Term 60 22.7 59 24.2 81 30.6
Persistent 150 56.8 69 28.3 90 34.0
Total 264 100.0 244 100.0 265 100.0
Table 3. User retention breakdown for Information Security scholars compared to the sub-fields that are similar in size.
4.3. Twitter-Mastodon Cross-platform Analysis
Since our attrition analysis showed that only a few users remained active on Mastodon, we conducted a cross-platform analysis to analyze their account state on Twitter. Approximately 50% of the scholars voluntarily provided their Twitter usernames, which were accurate and traceable to their Twitter profile pages. Using the Official Twitter Basic API, we collected the profile information of these accounts. Despite some account deletions, we successfully retrieved 3,131 Twitter accounts that could be matched to Mastodon accounts. Next, we retrieved the most recent posting times of these Twitter accounts to assess their activity status.
Based on their posting behavior, we categorized the Twitter account activity into the following statuses:
Locked: Scholar’s account is locked down, meaning their posts are private and cannot be accessed.
Inactive: Scholar’s latest tweet was posted before December 31st, 2022, and their posts were made before this date, indicating no recent activity since the start of the year.
Semi-active: Scholar’s latest tweet was posted between December 31st, 2022 and November 1, 2023. This indicates some activity during the year but no recent posts after November, 2023, the last timestamp of our Mastodon tracking.
Active: Scholar’s latest tweet was posted after November 1st, 2023, indicating continuous activity and engagement on the platform.
Using the above categorization, we calculated the count and connection statistics for each category, and all statistics are presented in Table 4 (left). We observe that 79.72% of scholars are still active on Twitter, indicating no signs of migration or intention to fully migrate. These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform. Some active users still make their Mastodon information accessible to the community. For instance, one scholar noted:
”Hi ya’ll for those moving over to #Mastodon you can currently find me ¡Scholar’s Mastodon Account¿.”
This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon. Furthermore, 7.03% of the scholars are semi-active users on Twitter. They did not stop using Twitter during the migration, but have reduced their presence gradually during a one-year grace period, potentially transitioning away from Twitter entirely.
In contrast, 2.91% of the scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter. For example, one scholar’s last post reads,
”I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!”
Approximately 10.35% of the scholars chose to lock their Twitter accounts, often signaling their departure from the platform by leaving a message in their profile directing followers to their Mastodon accounts.
Figure 6. Connection metrics mapping between mastodon and Twitter.
Count Following Count Follower Count One-time Short-term Long-term Persistent
Active 2496 Mean:1989.07, std:3087.68 Mean:13208.76 , std:108495.54 374 578 684 860
Semi-active 220 Mean:1137.56 , std:1263.87 Mean:4467.28, std:19152.90 39 42 43 96
Inactive 91 Mean:814.56 , std:1062.39 Mean:2401.46 , std:5203.50 12 19 21 39
Locked 324 Mean:1859.72 , std:2300.91 Mean:6347.10 , std:31583.98 51 51 82 140
Table 4. Engagement statistics of 3,131 scholars and mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4 (right). Our analysis reveals that 49.31% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users). Among those who are inactive on Twitter or have locked their accounts, 67.95% are long-term or persistent users of Mastodon. However, this group of dedicated Mastodon users represents only 9.02% of the total scholars included in the cross-platform analysis. This suggests that while there is a core group of loyal users who have fully migrated to Mastodon, they make up a relatively small portion of the overall academic community.
We also plotted the pairwise follower and following counts for the Mastodon and Twitter accounts of the same 3,131 scholars, as shown in Figure 6. The difference in connection scale is clear: while scholars who migrated to Mastodon made efforts to establish a new community, it is nowhere near the scale of connections they built over the years on Twitter. Additionally, we plotted the Twitter account creation date against the total number of tweets, displayed in Figure A3. The color of the points represents the total number of likes received by each account. A great proportion of scholars created their Twitter accounts between 2008 and 2012, indicating that the deep, long-standing networks, high levels of engagement, and rich posting histories developed on Twitter are not easily replicable or discarded. This entrenched presence highlights the difficulty of transitioning to a new platform like Mastodon without sacrificing years of accumulated influence and community ties.
4.3.1. Alternatives to Mastodon
In addition to Mastodon, other centralized alternatives to Twitter have emerged in recent years, and many academics explored these venues. Unlike Mastodon, these platforms provided familiar interfaces and ease of transitioning. Those present on these alternative platforms often added their account information in their Twitter account descriptions to make it easier for others to discover and connect with them on these new platforms. To understand the extent to which academics were exploring the alternatives,we analyzed their usernames and user descriptions for mentions of these platforms. Specifically, we focused on mentions of Bluesky [^5] https://bsky.app/ and Threads [^6] https://www.threads.net/ using keyword matching as well as user name pattern matching. For Mastodon and Bluesky, we also included scholars who use decentralized platform-style usernames as their Twitter usernames (e.g., ¡username¿@¡servername¿).
Figure 7 shows an UpSet plot illustrating the number of users who explicitly mention each platform or a combination of platforms. In addition to 626 scholars mentioning Mastodon, 308 and 69 users also reference BlueSky and Threads, respectively, with some scholars providing account information for multiple platforms. This indicates that while these scholars have accounts on Mastodon, only about 20% actively direct their Twitter connections to it. The trend highlights that Mastodon is not the sole option for scholars migrating away from Twitter and faces significant competition from other emerging platforms. This competition further complicates Mastodon’s ability to establish itself as the primary platform for academic discourse.
Figure 7. Mentions of Platforms in Username and Profile Description.
In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously. Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community. Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
4.4. Analysis of Tweets on Mastodon Migration
To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 3). We plot the daily frequency of these tweets and engagement counts over time aggregated by week in Figure 8. Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement. After the migration period, although conversations were less frequent, they received more attention (likes), indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
Conversations reflect difficulties and concerns during the transition. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
”I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am ¡scholar’s Mastodon account¿.”
”As you know, I’m over on Mastodon too. Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does.”
However, this initial surge in discussion was short-lived. The frequency of related tweets sharply declined.
Figure 8. Dynamic frequency of Twitter conversations related to academics migrating to Mastodon and corresponding engagement metrics (connection and like counts), aggregated by week.
Users expressed frustration over the platform’s lack of centralized theme and purpose, and its fragmented community structure.
”Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity. but the *big* servers that most ppl are on are… kinda academic? kinda techy? very… not fun? it feels like linkedin or facebook groups to a lot of people.”
”My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.”
Despite these challenges, ongoing efforts to promote Mastodon indicate a persistent belief in its potential. For instance, in June 2023, one scholar posted,
”Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd”
Amongst the discussions, we again observe concurrent mentions of other platforms such as Bluesky and Threads. These have diluted the migration as users explore various platforms for different purposes. For instance,
”I’m curious re categories of people I follow and where they go. IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?).”
”Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic”
From a qualitative perspective, these conversations reflect both optimism in Mastodon’s potential and the practical concerns academics face during the migration.
5. Discussion and Conclusion
In this work, we explored the migratory dynamics of Academic Twitter communities to Mastodon. Through a series of analyses, we highlight that building a community on a new platform requires more than a coordinated effort and availability of alternatives.
Results from longitudinal analysis of Mastodon user data showed that the initial surge in sign-ups did not translate into sustained long-term user engagement. Most recent user activity on Twitter reveals that many academics in our dataset returned to using Twitter actively. Further analysis of engagement metrics reveals that the level of established history, as well as the strong communities established on Twitter, with some over a decade, proved too significant to overcome. Without a complete migration of user history, and established community, professionals will naturally gravitate back to familiar platforms where their communities are already established through years.
Several factors likely contributed to the failure of Mastodon to retain users. First, its decentralized structure comes with a steep learning curve compared to Twitter. Furthermore, competition from more user-friendly platforms like Bluesky and Threads potentially drew users away from Mastodon. Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter. While some users expressed doubts, certain groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition.
One significant limitation of this study is the inconsistency in data collection timing across different platforms. This discrepancy could potentially introduce biases in the analysis, as activity levels of scholars might vary depending on when the data was captured. Additionally, incorrect or incomplete social media handles provided by the scholars led to gaps in data that we could not manually resolve due to the scale of the dataset. Moreover, the changes to platform access (such as Twitter API rate limitations) restricted our ability to gather sufficient data throughout the study. However, we believe our findings from the remaining data is substantial and provides valuable insights for the cross platform analyses.
Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement. While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions. However, the Academic Twitter community’s large scale attempt to migrate from Twitter demonstrates their adaptability and resilience. If any group could successfully make this migration, it would be a community like Academic Twitter, united by a shared purpose and strong networks. Yet, as encouraging as these efforts are, there is a long way to go.